# On the optimality of coin-betting for mean estimation

### v1

@@ Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This letter establishes that such coin-betting formulation is optimal among all possible algorithmic frameworks for constructing confidence sequences that build on e-variables and sequential hypothesis testing.
keywords:
mean estimation , confidence sequences , coin-betting , e-variables , sequential hypothesis testing
journal: Statistics & Probability Letters
@@ 1 Introduction
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022, 2023). E-variables, namely non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
The purpose of this letter is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of an e-variable-based algorithmic framework can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such algorithmic approach. The main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from the existing literature, which is primarily focused on the optimality of an individual e-variable (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024) or of a single wealth super-martingale (e.g., admissibility in Ramdas et al. 2022). The author hopes that this new viewpoint will inspire further research.
@@
Endow the interval $[0,1]$ with the standard topology, and let $\mathcal{P}$ be the set of all Borel probability measures on $[0,1]$. For $x$ in $[0,1]$, $\delta_{x}$ denotes the Dirac unit mass on $x$. For $P\in\mathcal{P}$ and a Borel measurable function $f$ on $[0,1]$, $\langle P,f(X)\rangle$ (or more compactly $\langle P,f\rangle$) denotes the expectation of $f$ under $X\sim P$. Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$). For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d. draws from $P\in\mathcal{P}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
@@ 2 Algorithmic mean estimation
Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
@@ 2.1 Coin-betting confidence sequence
Definition 1 (Coin-betting game)
@@ 2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. We call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}$. An e-variable (with respect to $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\langle P,E\rangle\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
@@
Definition 2 (Testing game)
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}$ and an e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$, a player
  • 1.
    picks [^2] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • 2.
    observes a new data-point $x_{t}\in[0,1]$;
  • 3.
    gets a reward $\log E_{t}(x_{t})$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing game.
The above game is at the core of e-variable testing (see, e.g., Ramdas and Wang 2024 for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1
Let $\mathcal{H}\subseteq\mathcal{P}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
@@
Proof 1
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$$\qed$
2.3 Mean estimation via e-variables
We can leverage the testing game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle=\mu\}\,,$ (1)
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Lemma 2
@@ 2.3 Mean estimation via e-variables
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\langle P,E_{t}\rangle=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
@@
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, the answer turns out to be negative, as we will demonstrate throughout the remainder of this letter.
@@ 3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on $[0,1]$. Given two functions $f,f^{\prime}:[0,1]\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in[0,1]$. If $f\succeq f^{\prime}$ and there exists a point $x\in[0,1]$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3
@@
Definition 4
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 3
@@
Proof 3
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$.
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing game of Definition 2.
Definition 5
If a majorising e-class is contained in every other majorising e-class, it is called optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 4
@@
Now, the final statements are a trivial consequence of what shown above and Lemma 3$\qed$
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, the optimal approach to obtain confidence sequences consists in restricting the testing game to the optimal e-class, for each mean candidate value $\mu$. We will show next that this coincides precisely with the coin-betting formulation.

### v2

@@ Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for the mean of bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This work discusses the optimality of such coin-betting formulation among algorithmic frameworks building on e-variables methods to test and estimate the mean of bounded random variables.
@@ 1 Introduction
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023). E-variables, non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice can depend on the past observations. Our main goal is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of such round-wise testing-by-betting game can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such betting approach. One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from much of the existing literature, which is primarily focused on the optimality of an individual e-variable or wealth process (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024, or admissibility in Ramdas et al. 2022a).
An alternative way to frame the first part of this work is as a characterisation of the family of single-round e-variables for testing the mean of a bounded random variable. In a similar spirit, since the appearance of the first version of this manuscript, two works have derived more general results for testing hypotheses defined by linear constraints (Clerico, 2024; Larsson et al., 2025), of which the present setting is a particular case. However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
We remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024). These allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d. or fixed conditional mean). In such cases, restricting the player to select a single-round e-variable at each step can be limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b). The second part of this work considers this broader setting. For technical convenience, we only focus on variables with finite discrete support, though we conjecture our conclusions hold more generally. We establish that when the sequence has fixed conditional mean, coin betting remains optimal even among testing methods based on multi-round e-variables and e-processes. However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d. observations.
@@
We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$. For $z\in\mathcal{Z}$, $\delta_{z}$ denotes the Dirac unit mass on $z$. For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$. Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ denotes the conditional expectation. Typically, $\mathcal{G}$ will be the sigma-field generated by some random variable $X$. In such case, we will write $\mathbb{E}_{P}[Z|X]$.
We will also consider measures on product spaces. For $T\geq 1$, we write $\mathcal{P}_{\mathcal{Z}^{T}}$ the set of Borel probability measure on $\mathcal{Z}^{T}$ (endowed with the product topology). We will used $\otimes$ to denote the direct product of measures. For instance, given $P$ and $Q$ in $\mathcal{P}_{\mathcal{Z}}$, $P\otimes Q$ will be the element of $\mathcal{P}_{Z^{2}}$ that encode the law of $(Z_{1},Z_{2})$, where $Z_{1}\sim P$ and $Z_{2}\sim Q$ are independent.
For any two given integers $s$ and $t$, with $s\leq t$, $[s:t]$ denotes the set of integers between $s$ and $t$ (included). For an integer $T$, given a vector $(x_{1},\dots,x_{T})$, we often represent it compactly as $x^{T}$ (upper indices). At times, we will also use the notation $x^{t:T}$ (with $t\in[1:T]$), to denote the vector $(x_{t},x_{t+1},\dots,x_{T})$. Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$). Sometimes we will also use the notation $s^{T_{0}:\infty}$ to denote $(s_{t})_{t\geq T_{0}}$, or simply $s^{\infty}$ if $T_{0}$ is clear from the context. For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d. draws from $P\in\mathcal{P}_{\mathcal{X}}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
@@ 2 Algorithmic mean estimation
Fix a Borel set $\mathcal{X}\subseteq\mathbb{R}$. Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
@@
We stress here that, for the time being, we assume that the sequence $(X_{t})_{t\geq 1}$ is i.i.d., and we shall not challenge this assumption, regardless of whether or not the observed data appear to exhibit such behaviour. In a way, this perspective aligns naturally with the goal of constructing confidence sequences, as simultaneously testing for the mean and the i.i.d. assumption might lead to rejecting everything and returning empty confidence sets because “the data do not look i.i.d.”. Under the i.i.d. model, mean estimation is well posed: we assume the existence of a fixed distribution $P^{\star}$ independently generating each observation, and our task is to estimate its mean. We remark that the independence assumption could be relaxed by only asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$. The arguments we present next for the i.i.d. setting carry over with essentially no change to this less restrictive case. However, to keep the exposition clearer, we focus here solely on the independent case. The conditional setting will be addressed in the second part of this work, where we consider the more general scenario in which dependencies between observations are not assumed but are instead challenged by the statistical test.
@@ 2.1 Coin-betting confidence sequence
Definition 1 (Coin-betting game).
@@ 2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. Here, we call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}_{[0,1]}$. An e-variable (for $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
@@
Definition 2 (Testing-by-betting game).
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$,
  • the player picks [^3] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • the player observes a new data-point $x_{t}\in[0,1]$;
  • the player earns a reward $\log E_{t}(x_{t})$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing-by-betting game.
The above game is an instance of e-variable testing (see Ramdas and Wang, 2024, for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1.
Let $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
@@
Proof.
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$. ∎
2.3 Mean estimation via sequential betting
We can leverage the testing-by-betting game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Proposition 1.
@@ 2.3 Mean estimation via sequential betting
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. ∎
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\mathbb{E}_{P}[E_{t}]=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
@@
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, for the round-wise testing-by-betting framework that we are considering the answer turns out to be negative.
@@ 3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on a set $\mathcal{X}$. Given two functions $f,f^{\prime}:\mathcal{X}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and there exists a point $x\in\mathcal{X}$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3.
@@
Definition 4.
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}^{\prime}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 2.
@@
Proof.
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$. ∎
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing-by-betting game of Definition 2.
Definition 5.
If a majorising e-class is contained in every other majorising e-class, we call it optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. [^4] We refer the reader to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class. Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on $[0,1]$) is a hypothesis for which the set of maximal e-variables is not a majorising e-class. Hence, $\mathcal{H}$ does not admit an optimal e-class. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 3.
@@
Now, the final statements are a trivial consequence of what was shown above and Lemma 2. ∎
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing-by-betting game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to obtain confidence sequences consists in restricting game to the optimal e-class, for each mean candidate value $\mu$. We show next that this coincides precisely with the coin-betting formulation.